OpenStudy (anonymous):

write the equation of the line parallel to the line 4x-6y=-1 and contains the x intercept of 3x-2y=12

OpenStudy (anonymous):

slope = 2/3 , x-intercept = (4.0) ===> equation is y =(2/3) x+b ==> b=-8/3 ===> y=(2/3)x-(8/3)

OpenStudy (anonymous):

Okay, first off, we need a line with the same slope as 4x-6y=-1 In slope-intercept form this becomes: 6y=4x+1 or y=(2/3)x+(1/6) So our slope is 2/3 Now we also need an x-intercept which is the same as the x-intercept of the line 3x-2y=12 Setting y=0, we solve for the x-int: 3x-2(0)=12 3x=12 x=4 So our x-int is the point (4,0). We now have a point (4,0) and a slope (=2/3) which is all we need to describe a line. Point-slope form gives us our line: y-0=(2/3)(x-4) Rearranging, we get: y=(2/3)x-(8/3)

OpenStudy (anonymous):

agreed.

OpenStudy (anonymous):

So, first thing's first: bust out your list of trig identities and find: \[1+\tan ^2x\]